### v2

@@ B.2.3 With noise at level 10%
B.2.4 Real application: maximizing the silhouette score
| Gini Generalized | 1.18 | 0.4915 | 0.4161 | 4.88 | 0.6339 | 0.6363 | 2.48 | 0.6252 | 0.5650 | 1.72 | 0.6186 | 0.3495 | 1.84 | 0.8214 | 0.5520 | 2.14 | 0.4772 | 0.3409 | 2.36 | 0.4466 | 0.1273 | 1.62 | 0.7052 | 0.0759 |
| Gini G. silhouette | 4.62 | 0.3410 | 0.3080 | 4.60 | 0.6247 | 0.6275 | 2.20 | 0.4610 | 0.5560 | 1.76 | 0.5581 | 0.3132 | 1.88 | 0.6273 | 0.5312 | 2.34 | 0.4198 | 0.3443 | 2.82 | 0.2689 | 0.0295 | 1.10 | 0.6991 | 0.0432 |
| Gini | - | 0.3897 | 0.3535 | - | 0.5822 | 0.5816 | - | 0.4645 | 0.5360 | - | 0.4916 | 0.3330 | - | 0.7131 | 0.5408 | - | 0.4082 | 0.3475 | - | 0.1467 | 0.1273 | - | 0.6828 | 0.1174 |
| Euclidean | - | 0.3147 | 0.2882 | - | 0.6077 | 0.6115 | - | 0.4123 | 0.5130 | - | 0.4239 | 0.2672 | - | 0.6977 | 0.5296 | - | 0.3183 | 0.3213 | - | 0.5445 | 0.1227 | - | 0.4288 | 0.0155 |
| Manhattan | - | 0.3586 | 0.3280 | - | 0.6083 | 0.6123 | - | 0.4123 | 0.5130 | - | 0.4113 | 0.2573 | - | 0.7204 | 0.5728 | - | 0.3247 | 0.3443 | - | 0.5992 | 0.0977 | - | 0.4288 | 0.0155 |
| Minkowski | 3.0 | 0.3200 | 0.2910 | 3.0 | 0.6085 | 0.6123 | 3.0 | 0.4123 | 0.5130 | 3.0 | 0.4254 | 0.2605 | 3.0 | 0.6475 | 0.4928 | 3.0 | 0.3208 | 0.2909 | 3.0 | 0.5315 | 0.1386 | 3.0 | 0.4288 | 0.0155 |
| Cosine | - | 0.4013 | 0.3621 | - | 0.5974 | 0.6006 | - | 0.3084 | 0.4840 | - | 0.5479 | 0.2934 | - | 0.7292 | 0.5088 | - | 0.3259 | 0.3770 | - | 0.5952 | 0.4886 | - | 0.6376 | 0.1575 |
| Canberra | - | 0.5419 | 0.4505 | - | 0.5901 | 0.5948 | - | 0.5435 | 0.5250 | - | 0.5651 | 0.2013 | - | 0.4913 | 0.3744 | - | 0.6109 | 0.4388 | - | 0.2348 | 0.1136 | - | 0.6864 | 0.1798 |
| Hellinger | - | 0.5754 | 0.5021 | - | 0.6015 | 0.5868 | - | 0.3364 | 0.4740 | - | 0.4159 | 0.1985 | - | 0.5745 | 0.3760 | - | 0.4612 | 0.4304 | - | 0.3319 | 0.2273 | - | 0.7150 | 0.0414 |
| Jensen-Shannon | - | 0.1289 | 0.3590 | - | 0.3085 | 0.5554 | - | 0.3095 | 0.5050 | - | 0.2548 | 0.0363 | - | 0.4687 | 0.1808 | - | 0.2361 | 0.0230 | - | 0.2621 | 0.1523 | - | 0.7150 | 0.0588 |
| PearsonChi2 | - | 0.1289 | 0.3590 | - | 0.5043 | 0.5094 | - | 0.2789 | 0.4960 | - | 0.4802 | 0.2644 | - | 0.4578 | 0.3616 | - | 0.5301 | 0.4550 | - | 0.5397 | 0.4955 | - | 0.6623 | 0.1678 |
| VicisSymmetric1 | - | 0.5412 | 0.5070 | - | 0.5225 | 0.5109 | - | 0.2954 | 0.4970 | - | 0.3202 | 0.1551 | - | 0.4240 | 0.3840 | - | 0.5324 | 0.4189 | - | 0.3484 | 0.1250 | - | 0.5064 | 0.4039 |
| Hassanat | - | 0.4226 | 0.3821 | - | 0.5556 | 0.5517 | - | 0.5754 | 0.5750 | - | 0.5199 | 0.4033 | - | 0.5938 | 0.4464 | - | 0.5128 | 0.4769 | - | 0.3946 | 0.3250 | - | 0.6958 | 0.1400 |
| Gini Generalized | 3 3.52 | 0.0998 | 0.1 | 3 4.10 | 0.1913 | 0.1738 | 2 2.70 | 0.9234 | 0.9157 | 2 1.22 | 0.6339 | 0.5956 | 3 3.12 | 0.3741 | 0.2978 | 3 4.52 | 0.2946 | 0.2801 | 2 1.66 | 0.7693 | 0.6145 | 6 2.54 | 0.1492 | 0.0699 |
| Gini G. Silhouette | 3 1.40 | 0.08 | 0.1 | 3 1.72 | 0.0902 | 0.0890 | 2 1.12 | 0.8840 | 0.8577 | 2 2.38 | 0.5746 | 0.5423 | 3 1.88 | 0.3485 | 0.2670 | 3 1.10 | 0.2207 | 0.2186 | 2 1.10 | 0.7609 | 0.5797 | 6 1.14 | 0.0798 | 0.0420 |
| Gini | 3 2 | 0.0828 | 0.1 | 3 2 | 0.1172 | 0.1005 | 2 2 | 0.9129 | 0.9052 | 2 2 | 0.5985 | 0.5618 | 3 2 | 0.3419 | 0.2644 | 3 2 | 0.2596 | 0.2517 | 2 2 | 0.6891 | 0.6101 | 6 2 | 0.0584 | 0.0327 |
| Euclidean | 3 - | 0.0757 | 0.0933 | 3 - | 0.0610 | 0.0502 | 2 - | 0.8802 | 0.8524 | 2 - | 0.5636 | 0.5373 | 3 - | 0.3520 | 0.2349 | 3 - | 0.2088 | 0.2269 | 2 - | 0.4873 | 0.5590 | 6 - | 0.0388 | 0.0186 |
| Manhattan | 3 - | 0.0788 | 0.1 | 3 - | 0.0680 | 0.0613 | 2 - | 0.8782 | 0.8489 | 2 - | 0.5818 | 0.5518 | 3 - | 0.3531 | 0.2567 | 3 - | 0.2057 | 0.2234 | 2 - | 0.4873 | 0.5590 | 6 - | 0.0628 | 0.0279 |
| Minkowski | 3 3 | 0.0774 | 0.0933 | 3 3 | 0.0643 | 0.0502 | 2 3 | 0.8812 | 0.8542 | 2 3 | 0.5784 | 0.5419 | 3 3 | 0.3498 | 0.2311 | 3 3 | 0.2100 | 0.2281 | 2 3 | 0.4873 | 0.5590 | 6 3 | 0.0359 | 0.0186 |
| Cosine | 3 - | 0.0188 | 0.02 | 3 - | 0.1370 | 0.1119 | 2 - | 0.9128 | 0.9051 | 2 - | 0.5627 | 0.5325 | 3 - | 0.3945 | 0.3055 | 3 - | 0.1984 | 0.2186 | 2 - | 0.6770 | 0.6725 | 6 - | 0.0599 | 0.0280 |
| Canberra | 3 - | 0.0386 | 0.0467 | 3 - | 0.3990 | 0.3430 | 2 - | 0.9237 | 0.9227 | 2 - | 0.5407 | 0.5225 | 3 - | 0.4449 | 0.3979 | 3 - | 0.2292 | 0.2234 | 2 - | 0.5208 | 0.5203 | 6 - | 0.1123 | 0.0421 |
| Hellinger | 3 - | 0.2906 | 0.2933 | 3 - | 0.2646 | 0.2740 | 2 - | 0.7313 | 0.7156 | 2 - | 0.3424 | 0.4901 | 3 - | 0.5476 | 0.3854 | 3 - | 0.3701 | 0.3275 | 2 - | 0.6701 | 0.5725 | 6 - | 0.1579 | 0.1728 |
| JensenShannon | 3 - | 0.4642 | 0.46 | 3 - | 0.0543 | 0.1352 | 2 - | 0.1272 | 0.3268 | 2 - | 0.5549 | 0.5531 | 3 - | 0.5241 | 0.5378 | 3 - | 0.3695 | 0.4196 | 2 - | 0.6830 | 0.6319 | 6 - | 0.0 | 0.0 |
| PearsonChi2 | 3 - | 0.2688 | 0.2867 | 3 - | 0.4505 | 0.4865 | 2 - | 0.8890 | 0.8683 | 2 - | 0.5973 | 0.5823 | 3 - | 0.3998 | 0.2990 | 3 - | 0.3644 | 0.3509 | 2 - | 0.6705 | 0.5913 | 6 - | 0.2025 | 0.1457 |
| VicisSymmetric1 | 3 - | 0.0667 | 0.0333 | 3 - | 0.2341 | 0.2037 | 2 - | 0.8656 | 0.8420 | 2 - | 0.4488 | 0.5296 | 3 - | 0.4766 | 0.2786 | 3 - | 0.4819 | 0.4031 | 2 - | 0.6502 | 0.5652 | 6 - | 0.1515 | 0.1260 |
| Hassanat | 3 - | 0.0497 | 0.06 | 3 - | 0.3990 | 0.3598 | 2 - | 0.9240 | 0.9157 | 2 - | 0.5834 | 0.5470 | 3 - | 0.3772 | 0.2490 | 3 - | 0.3563 | 0.2931 | 2 - | 0.7858 | 0.7739 | 6 - | 0.0553 | 0.0279 |
@@ B.2.4 Real application: maximizing the silhouette score
B.3 Additional experiments on agglomerative hierarchical clustering with 10% level of noise
| Gini Generalized | 3 1.10 | 0.4114 | 0.6410 | 3 5.07 | 0.8009 | 0.6888 | 3 1.38 | 0.5402 | 0.5380 | 3 2.78 | 0.5563 | 0.5974 | 3 1.35 | 0.6726 | 0.6063 | 3 1.65 | 0.7003 | 0.7353 | 3 3.76 | 0.8622 | 0.8182 | 3 1.10 | 0.5132 | 0.7150 |
| Gini | 3 | 0.4114 | 0.6410 | 3 | 0.3885 | 0.5503 | 3 | 0.5117 | 0.5070 | 3 | 0.4324 | 0.5380 | 3 | 0.5973 | 0.6416 | 3 | 0.5409 | 0.7353 | 3 | 0.4582 | 0.6727 | 3 | 0.5132 | 0.7150 |
| Euclidean | 3 | 0.4114 | 0.6410 | 3 | 0.7538 | 0.6618 | 3 | 0.3190 | 0.4850 | 3 | 0.3424 | 0.5446 | 3 | 0.5566 | 0.5840 | 3 | 0.5409 | 0.7353 | 3 | 0.4592 | 0.6773 | 3 | 0.5132 | 0.7150 |
| Euclidean Ward | 3 | 0.6712 | 0.6667 | 3 | 0.7137 | 0.6550 | 3 | 0.5420 | 0.5435 | 3 | 0.3074 | 0.5545 | 3 | 0.7044 | 0.6298 | 3 | 0.5699 | 0.7549 | 3 | 0.8195 | 0.7891 | 3 | 0.5880 | 0.7668 |
| Manhattan | 3 | 0.5322 | 0.6439 | 3 | 0.7573 | 0.6720 | 3 | 0.4272 | 0.5110 | 3 | 0.3424 | 0.5446 | 3 | 0.6252 | 0.6336 | 3 | 0.5409 | 0.7353 | 3 | 0.4592 | 0.6773 | 3 | 0.5132 | 0.7150 |
| Minkowski | 3 3 | 0.4114 | 0.6410 | 3 3 | 0.7385 | 0.6684 | 3 3 | 0.4131 | 0.4920 | 3 3 | 0.3424 | 0.5446 | 3 3 | 0.5247 | 0.5649 | 3 3 | 0.5409 | 0.7353 | 3 3 | 0.4592 | 0.6773 | 3 3 | 0.5132 | 0.7150 |
| Cosine | 3 | 0.4114 | 0.6410 | 3 | 0.6432 | 0.6443 | 3 | 0.4021 | 0.4770 | 3 | 0.3630 | 0.5479 | 3 | 0.6220 | 0.6160 | 3 | 0.5409 | 0.7353 | 3 | 0.5899 | 0.7249 | 3 | 0.5132 | 0.7150 |
| Canberra | 3 | 0.6065 | 0.6496 | 3 | 0.6126 | 0.5685 | 3 | 0.2277 | 0.4770 | 3 | 0.5402 | 0.6502 | 3 | 0.6541 | 0.6335 | 3 | 0.5409 | 0.7353 | 3 | 0.4592 | 0.6773 | 3 | 0.5132 | 0.7150 |
| Hellinger | 3 | 0.4114 | 0.6410 | 3 | 0.7139 | 0.6553 | 3 | 0.3967 | 0.4770 | 3 | 0.3647 | 0.5479 | 3 | 0.6282 | 0.6385 | 3 | 0.5409 | 0.7353 | 3 | 0.5768 | 0.6818 | 3 | 0.5132 | 0.7150 |
| JensenShannon | 3 | 0.4114 | 0.6410 | 3 | 0.3089 | 0.5554 | 3 | 0.2277 | 0.4770 | 3 | 0.3649 | 0.5479 | 3 | 0.6515 | 0.6000 | 3 | 0.5409 | 0.7353 | 3 | 0.5783 | 0.7158 | 3 | 0.5132 | 0.7150 |
| PearsonChi2 | 3 | 0.6615 | 0.6467 | 3 | 0.5251 | 0.5401 | 3 | 0.3848 | 0.4680 | 3 | 0.3424 | 0.5446 | 3 | 0.5378 | 0.5760 | 3 | 0.6036 | 0.7418 | 3 | 0.5377 | 0.6795 | 3 | 0.5132 | 0.7150 |
| VicisSymmetric1 | 3 | 0.6579 | 0.6581 | 3 | 0.5517 | 0.5816 | 3 | 0.3188 | 0.4780 | 3 | 0.4889 | 0.5446 | 3 | 0.6336 | 0.6096 | 3 | 0.5409 | 0.7353 | 3 | 0.4592 | 0.6773 | 3 | 0.5132 | 0.7150 |
| Hassanat | 3 | 0.7739 | 0.6496 | 3 | 0.4520 | 0.5904 | 3 | 0.3807 | 0.4820 | 3 | 0.6616 | 0.7096 | 3 | 0.6426 | 0.5920 | 3 | 0.5409 | 0.7353 | 3 | 0.4592 | 0.6773 | 3 | 0.5132 | 0.7150 |

### v3

@@ 6 Conclusion
Real application: maximizing the silhouette score
Ionosphere Banknote German Heart Balance Haberman Wholesale Indian Liver Patient Distance #P #R #P #R #P #R #P #R #P #R #P #R #P #R #P #R Gini Generalized 1.18 0.4915 0.4161 4.88 0.6339 0.6363 2.48 0.6252 0.5650 1.72 0.6186 0.3495 1.84 0.8214 0.5520 2.14 0.4772 0.3409 2.36 0.4466 0.1273 1.62 0.7052 0.0759 Gini G. silhouette 4.62 0.3410 0.3080 4.60 0.6247 0.6275 2.20 0.4610 0.5560 1.76 0.5581 0.3132 1.88 0.6273 0.5312 2.34 0.4198 0.3443 2.82 0.2689 0.0295 1.10 0.6991 0.0432 Gini - 0.3897 0.3535 - 0.5822 0.5816 - 0.4645 0.5360 - 0.4916 0.3330 - 0.7131 0.5408 - 0.4082 0.3475 - 0.1467 0.1273 - 0.6828 0.1174 Euclidean - 0.3147 0.2882 - 0.6077 0.6115 - 0.4123 0.5130 - 0.4239 0.2672 - 0.6977 0.5296 - 0.3183 0.3213 - 0.5445 0.1227 - 0.4288 0.0155 Manhattan - 0.3586 0.3280 - 0.6083 0.6123 - 0.4123 0.5130 - 0.4113 0.2573 - 0.7204 0.5728 - 0.3247 0.3443 - 0.5992 0.0977 - 0.4288 0.0155 Minkowski 3.0 0.3200 0.2910 3.0 0.6085 0.6123 3.0 0.4123 0.5130 3.0 0.4254 0.2605 3.0 0.6475 0.4928 3.0 0.3208 0.2909 3.0 0.5315 0.1386 3.0 0.4288 0.0155 Cosine - 0.4013 0.3621 - 0.5974 0.6006 - 0.3084 0.4840 - 0.5479 0.2934 - 0.7292 0.5088 - 0.3259 0.3770 - 0.5952 0.4886 - 0.6376 0.1575 Canberra - 0.5419 0.4505 - 0.5901 0.5948 - 0.5435 0.5250 - 0.5651 0.2013 - 0.4913 0.3744 - 0.6109 0.4388 - 0.2348 0.1136 - 0.6864 0.1798 Hellinger - 0.5754 0.5021 - 0.6015 0.5868 - 0.3364 0.4740 - 0.4159 0.1985 - 0.5745 0.3760 - 0.4612 0.4304 - 0.3319 0.2273 - 0.7150 0.0414 Jensen-Shannon - 0.1289 0.3590 - 0.3085 0.5554 - 0.3095 0.5050 - 0.2548 0.0363 - 0.4687 0.1808 - 0.2361 0.0230 - 0.2621 0.1523 - 0.7150 0.0588 PearsonChi2 - 0.1289 0.3590 - 0.5043 0.5094 - 0.2789 0.4960 - 0.4802 0.2644 - 0.4578 0.3616 - 0.5301 0.4550 - 0.5397 0.4955 - 0.6623 0.1678 VicisSymmetric1 - 0.5412 0.5070 - 0.5225 0.5109 - 0.2954 0.4970 - 0.3202 0.1551 - 0.4240 0.3840 - 0.5324 0.4189 - 0.3484 0.1250 - 0.5064 0.4039 Hassanat - 0.4226 0.3821 - 0.5556 0.5517 - 0.5754 0.5750 - 0.5199 0.4033 - 0.5938 0.4464 - 0.5128 0.4769 - 0.3946 0.3250 - 0.6958 0.1400
Iris Wine Breast-Cancer Sonar QSAR Vehicle Australian Glass Distance #P #R #P #R #P #R #P #R #P #R #P #R #P #R #P #R Gini Generalized 3 3.52 0.0998 0.1 3 4.10 0.1913 0.1738 2 2.70 0.9234 0.9157 2 1.22 0.6339 0.5956 3 3.12 0.3741 0.2978 3 4.52 0.2946 0.2801 2 1.66 0.7693 0.6145 6 2.54 0.1492 0.0699 Gini G. Silhouette 3 1.40 0.08 0.1 3 1.72 0.0902 0.0890 2 1.12 0.8840 0.8577 2 2.38 0.5746 0.5423 3 1.88 0.3485 0.2670 3 1.10 0.2207 0.2186 2 1.10 0.7609 0.5797 6 1.14 0.0798 0.0420 Gini 3 2 0.0828 0.1 3 2 0.1172 0.1005 2 2 0.9129 0.9052 2 2 0.5985 0.5618 3 2 0.3419 0.2644 3 2 0.2596 0.2517 2 2 0.6891 0.6101 6 2 0.0584 0.0327 Euclidean 3 - 0.0757 0.0933 3 - 0.0610 0.0502 2 - 0.8802 0.8524 2 - 0.5636 0.5373 3 - 0.3520 0.2349 3 - 0.2088 0.2269 2 - 0.4873 0.5590 6 - 0.0388 0.0186 Manhattan 3 - 0.0788 0.1 3 - 0.0680 0.0613 2 - 0.8782 0.8489 2 - 0.5818 0.5518 3 - 0.3531 0.2567 3 - 0.2057 0.2234 2 - 0.4873 0.5590 6 - 0.0628 0.0279 Minkowski 3 3 0.0774 0.0933 3 3 0.0643 0.0502 2 3 0.8812 0.8542 2 3 0.5784 0.5419 3 3 0.3498 0.2311 3 3 0.2100 0.2281 2 3 0.4873 0.5590 6 3 0.0359 0.0186 Cosine 3 - 0.0188 0.02 3 - 0.1370 0.1119 2 - 0.9128 0.9051 2 - 0.5627 0.5325 3 - 0.3945 0.3055 3 - 0.1984 0.2186 2 - 0.6770 0.6725 6 - 0.0599 0.0280 Canberra 3 - 0.0386 0.0467 3 - 0.3990 0.3430 2 - 0.9237 0.9227 2 - 0.5407 0.5225 3 - 0.4449 0.3979 3 - 0.2292 0.2234 2 - 0.5208 0.5203 6 - 0.1123 0.0421 Hellinger 3 - 0.2906 0.2933 3 - 0.2646 0.2740 2 - 0.7313 0.7156 2 - 0.3424 0.4901 3 - 0.5476 0.3854 3 - 0.3701 0.3275 2 - 0.6701 0.5725 6 - 0.1579 0.1728 JensenShannon 3 - 0.4642 0.46 3 - 0.0543 0.1352 2 - 0.1272 0.3268 2 - 0.5549 0.5531 3 - 0.5241 0.5378 3 - 0.3695 0.4196 2 - 0.6830 0.6319 6 - 0.0 0.0 PearsonChi2 3 - 0.2688 0.2867 3 - 0.4505 0.4865 2 - 0.8890 0.8683 2 - 0.5973 0.5823 3 - 0.3998 0.2990 3 - 0.3644 0.3509 2 - 0.6705 0.5913 6 - 0.2025 0.1457 VicisSymmetric1 3 - 0.0667 0.0333 3 - 0.2341 0.2037 2 - 0.8656 0.8420 2 - 0.4488 0.5296 3 - 0.4766 0.2786 3 - 0.4819 0.4031 2 - 0.6502 0.5652 6 - 0.1515 0.1260 Hassanat 3 - 0.0497 0.06 3 - 0.3990 0.3598 2 - 0.9240 0.9157 2 - 0.5834 0.5470 3 - 0.3772 0.2490 3 - 0.3563 0.2931 2 - 0.7858 0.7739 6 - 0.0553 0.0279
@@
Additional experiments on agglomerative hierarchical clustering with 10% level of noise
Ionosphere Banknote German Heart Balance Haberman Wholesale Indian Liver Patient Distance #P #R #P #R #P #R #P #R #P #R #P #R #P #R #P #R Gini Generalized 3 1.10 0.4114 0.6410 3 5.07 0.8009 0.6888 3 1.38 0.5402 0.5380 3 2.78 0.5563 0.5974 3 1.35 0.6726 0.6063 3 1.65 0.7003 0.7353 3 3.76 0.8622 0.8182 3 1.10 0.5132 0.7150 Gini 3 0.4114 0.6410 3 0.3885 0.5503 3 0.5117 0.5070 3 0.4324 0.5380 3 0.5973 0.6416 3 0.5409 0.7353 3 0.4582 0.6727 3 0.5132 0.7150 Euclidean 3 0.4114 0.6410 3 0.7538 0.6618 3 0.3190 0.4850 3 0.3424 0.5446 3 0.5566 0.5840 3 0.5409 0.7353 3 0.4592 0.6773 3 0.5132 0.7150 Euclidean Ward 3 0.6712 0.6667 3 0.7137 0.6550 3 0.5420 0.5435 3 0.3074 0.5545 3 0.7044 0.6298 3 0.5699 0.7549 3 0.8195 0.7891 3 0.5880 0.7668 Manhattan 3 0.5322 0.6439 3 0.7573 0.6720 3 0.4272 0.5110 3 0.3424 0.5446 3 0.6252 0.6336 3 0.5409 0.7353 3 0.4592 0.6773 3 0.5132 0.7150 Minkowski 3 3 0.4114 0.6410 3 3 0.7385 0.6684 3 3 0.4131 0.4920 3 3 0.3424 0.5446 3 3 0.5247 0.5649 3 3 0.5409 0.7353 3 3 0.4592 0.6773 3 3 0.5132 0.7150 Cosine 3 0.4114 0.6410 3 0.6432 0.6443 3 0.4021 0.4770 3 0.3630 0.5479 3 0.6220 0.6160 3 0.5409 0.7353 3 0.5899 0.7249 3 0.5132 0.7150 Canberra 3 0.6065 0.6496 3 0.6126 0.5685 3 0.2277 0.4770 3 0.5402 0.6502 3 0.6541 0.6335 3 0.5409 0.7353 3 0.4592 0.6773 3 0.5132 0.7150 Hellinger 3 0.4114 0.6410 3 0.7139 0.6553 3 0.3967 0.4770 3 0.3647 0.5479 3 0.6282 0.6385 3 0.5409 0.7353 3 0.5768 0.6818 3 0.5132 0.7150 JensenShannon 3 0.4114 0.6410 3 0.3089 0.5554 3 0.2277 0.4770 3 0.3649 0.5479 3 0.6515 0.6000 3 0.5409 0.7353 3 0.5783 0.7158 3 0.5132 0.7150 PearsonChi2 3 0.6615 0.6467 3 0.5251 0.5401 3 0.3848 0.4680 3 0.3424 0.5446 3 0.5378 0.5760 3 0.6036 0.7418 3 0.5377 0.6795 3 0.5132 0.7150 VicisSymmetric1 3 0.6579 0.6581 3 0.5517 0.5816 3 0.3188 0.4780 3 0.4889 0.5446 3 0.6336 0.6096 3 0.5409 0.7353 3 0.4592 0.6773 3 0.5132 0.7150 Hassanat 3 0.7739 0.6496 3 0.4520 0.5904 3 0.3807 0.4820 3 0.6616 0.7096 3 0.6426 0.5920 3 0.5409 0.7353 3 0.4592 0.6773 3 0.5132 0.7150
Iris Wine Breast Cancer Sonar Australian Glass QSAR Vehicle Distance #P #R #P #R #P #R #P #R #P #R #P #R #P #R #P #R Gini Generalized 3 2.37 0.8990 0.8867 3 4.27 0.7864 0.7187 2 2.50 0.9161 0.9050 2 2.54 0.4766 0.5336 3 2.22 0.7647 0.5957 6 1.20 0.5835 0.4485 3 1.36 0.6227 0.5982 4 4.95 0.6631 0.6028 Gini 3 - 0.8808 0.8533 3 - 0.5433 0.6575 2 - 0.8697 0.8312 2 - 0.3509 0.5239 3 - 0.7153 0.5826 6 - 0.5573 0.4765 3 - 0.3992 0.5738 4 - 0.4779 0.5035 Euclidean 3 - 0.8798 0.8133 3 - 0.5077 0.6345 2 - 0.7767 0.6520 2 - 0.4187 0.5239 3 - 0.4434 0.5565 6 - 0.4843 0.4068 3 - 0.5287 0.5982 4 - 0.6847 0.5330 Euclidean Ward 3 - 0.8773 0.8600 3 - 0.7941 0.7797 2 - 0.8872 0.8632 2 - 0.2947 0.5429 3 - 0.7664 0.6174 6 - 0.6302 0.4648 3 - 0.5360 0.5946 4 - 0.6592 0.5745 Manhattan 3 - 0.8945 0.8333 3 - 0.5104 0.6511 2 - 0.7767 0.6520 2 - 0.3741 0.5193 3 - 0.4434 0.5565 6 - 0.4024 0.4068 3 - 0.5540 0.5969 4 - 0.6878 0.5366 Minkowski 3 3 0.7308 0.6933 3 3 0.5104 0.6511 2 3 0.7767 0.6520 2 3 0.2848 0.5336 3 3 0.4434 0.5565 6 3 0.4319 0.4673 3 3 0.5657 0.6008 4 3 0.6028 0.5485 Cosine 3 - 0.6619 0.7667 3 - 0.4529 0.5901 2 - 0.7831 0.6660 2 - 0.2848 0.5336 3 - 0.4771 0.5580 6 - 0.4272 0.4626 3 - 0.3491 0.5905 4 - 0.4863 0.5047 Canberra 3 - 0.4870 0.6533 3 - 0.4728 0.6518 2 - 0.3940 0.6274 2 - 0.4386 0.5384 3 - 0.3095 0.5551 6 - 0.3332 0.4439 3 - 0.3491 0.5905 4 - 0.4879 0.4941 Hellinger 3 - 0.8781 0.7667 3 - 0.5793 0.6513 2 - 0.7755 0.6484 2 - 0.2848 0.5336 3 - 0.7562 0.5609 6 - 0.4082 0.4859 3 - 0.4302 0.6033 4 - 0.4774 0.5130 JensenShannon 3 - 0.4870 0.6533 3 - 0.2890 0.4878 2 - 0.3940 0.6274 2 - 0.2848 0.5336 3 - 0.5377 0.6101 6 - 0.1265 0.3552 3 - 0.3491 0.5905 4 - 0.2588 0.5071 PearsonChi2 3 - 0.8781 0.7667 3 - 0.6765 0.6964 2 - 0.6517 0.6379 2 - 0.3484 0.5143 3 - 0.7562 0.5609 6 - 0.4082 0.4859 3 - 0.3491 0.5905 4 - 0.4757 0.5213 VicisSymmetric1 3 - 0.4870 0.6533 3 - 0.3309 0.4045 2 - 0.3940 0.6274 2 - 0.3588 0.5193 3 - 0.5318 0.6014 6 - 0.4604 0.4623 3 - 0.3491 0.5905 4 - 0.2984 0.5024 Hassanat 3 - 0.7904 0.7867 3 - 0.6191 0.7365 2 - 0.8660 0.8314 2 - 0.3458 0.5288 3 - 0.6869 0.5768 6 - 0.5249 0.4907 3 - 0.3491 0.5905 4 - 0.4873 0.4917